Assessment 4

When testing for independence in a contingency table with 4 rows and 4 columns, there are _____________ degrees of freedom.
a.
6
b.
7
c.
9
d.
5
e.
4

Answer by math_tutor2020(3817)


r = number of rows = 4
c = number of columns = 4

df = degrees of freedom
df = (r-1)*(c-1)
df = (4-1)*(4-1)
df = 3*3
df = 9

Answer: There are 9 degrees of freedom